OpenStudy (anonymous):

State the horizontal asymptote of the rational function. f(x) = quantity x plus nine divided by quantity x squared plus two x plus three.

OpenStudy (anonymous):

\[f(x)=x+9/x ^{2}+2x+3\] is the equation

OpenStudy (anonymous):

If the degree of the first term in the numerator is lower than the degree of the first term of the denominator, then the horizontal asymptote is y=0. Example: \[f(x)=\frac{ x+2 }{ x^2-3x+2 }\] BUT If the degree of the first term in the numerator is higher than the degree of the first term of the denominator, then the horizontal asymptote is none. Example: \[f(x)=\frac{ x^2-3x+2 }{ x+10 }\]

OpenStudy (anonymous):

so none would be the answer?

OpenStudy (anonymous):

HOWEVER If the degree are the same in both numerator and denominator, you have to divide the coefficient of the first degree of the term of he numerator to the denominator's. Example: \[f(x)=\frac{ 2x^2-11 }{ x^2+9 }\] The horizontal slope here is y=2.

OpenStudy (anonymous):

Check my previous replies again.

OpenStudy (anonymous):

I already gave you an example. :D

OpenStudy (anonymous):

okay thank you